Nitrogen (II) oxide reacts with oxygen in the following reaction: Experiments were carried out to determine the orders of reaction with respect to NO and 02. The results of these experiments are shown below Experiment NO/mol-cm02l/ mol-cm Rate mol-cms 1.5 x 10 4.5 x 10 1.5 x 10 5 x 10 5 x 10 2.0 x 10 1 x10 1.9 x10 8.4 x10 a) Using the data, deduce the orders of reaction with respect to NO and O2 b) Write down the rate equation for this reaction and use it to calculate k. c) Explain the meaning of the term rate determining step? d) The Langmuir Adsorption Isotherm model assumes monolayer adsorption. There are two fundamental types of adsorption. Which kind of adsorption leads to monolayer adsorption? c) For the case in (d), sketch the surface coverage as a function of exposure to the adsorbate. f) What can cause deviations from ideal adsorption behavior? 8) Which isotherm would be more appropriate if the attraction were van der Waals type? How does this differ from the case in (d) in terms of how layers form on the surface?
Answer #1

a) To deduce the orders of the reaction, we need to compare the experiments.

To find order with respect to NO, we will compare experiment 1 and 2, because the [O2] is same in both.

rate1/rate2 = ([NO]1/[NO]2)x

(2.1*10-7)/(1.9*10-6)= ((1.5*10-5)/(4.5*10-5) )x

0.11 = (0.33)x

Taking log on both sides

log 0.11 = x log 0.33

x = 2

So, order with respect to [NO] is 2.

Similarly, to find order with respect to O2, we will compare experiment 1 and 3, because the [NO] is same in both.

rate1/rate3 = ([O2]1/[O2]3)y

(2.1*10-7)/(8.4*10-7)= ((0.5*10-5)/(2.0*10-5) )y

0.25 = (0.25)y

y = 1

So, order with respect to [O2] is 1.

b) Rate equation =>

rate = k [NO]2[O2]1

For experiment 1:

rate = 2.1*10-7 mol cm-3 s-1

[NO] = 1.5*10-5 mol cm-3

[O2] = 0.5*10-5 mol cm-3

Put all these values in rate equation

2.1*10-7 = k (1.5*10-5)2(0.5*10-5)

k = 1.87*108 cm3 mol-1 s-1

c) When a reaction is actually occuring through a number of steps, then the slowest step of that reaction is the rate determining step. This means that the rate of the whole reaction will depend on that step.

d) There are two fundamental types of adsorption: physisorption and chemisorption. Physisorption occurs due to vander waal forces and chemisorption occurs due to chemical bonds. Out of these two, chemisorption formed monolayer and physisorption forms multilayers.